# Complementary Probes of Warped Extra Dimension: Colliders, Gravitational Waves and Primordial Black Holes from Phase Transitions

### v1

@@ VI Collider and astrophysical data
The universal feature of 5D models is the presence of KK gravitons and scalar (radion) fluctuations propagating along the fifth dimension. As explained above, we are assuming the SM propagating on the TeV (or SM) brane and so the coupling of gravitons and the radion (localized toward the IR brane) depends on the distance between the SM brane and the IR brane. In such a scenario, the phenomenology is driven by the first graviton KK mode with mass $m_{h}\equiv\hat{m}_{h}\rho$ ($\hat{m}_{h}\simeq 3.83$) and the lightest radion state with mass $m_{\chi}\equiv\hat{m}_{\chi}\rho$ ($\hat{m}_{\chi}\simeq 0.01-0.1$), which have wave functions $h_{\mu\nu}(x)$ and $\chi(x)$, respectively. [^9] We are not considering the massless graviton zero mode which is very weakly coupled to the SM as its interaction is suppressed by $1/M_{P}$. These fields are coupled to the SM with Lagrangian Csaki et al. (2001)
@@
$\kappa_{h}$ $=\frac{2\sqrt{2}\pi}{N}\frac{1}{\rho}\,\left[\Theta\left(\frac{\rho-\rho_{T}}{\textrm{TeV}}\right)+\left(\frac{\rho}{\rho_{T}}\right)^{2}\frac{J_{2}(c_{1}\rho/\rho_{T})}{J_{2}(c_{1})}\left(\frac{\rho_{T}-\rho}{\textrm{TeV}}\right)\right]\,,$ (34)
@@
where $J_{n}(x)$ is a Bessel function of the first kind, and $\Theta(x)$ is the Heaviside step function equal to 1 (0) for $x\geq 0$ ($x<0$) which allows us to automatically swap from the high-energy to the low-energy setup when we cross the threshold $\rho=1\,$TeV. (As explained in Sec. II, we choose such a value of $\rho$ for concreteness, but other values of the order of 1 TeV would work as well.)
VI.1 Bounds from the KK graviton sector
The existence of the KK graviton resonances influence how the strength of the gravitational force scales with distance and how fast the star’s cores lose energy due to the additional emission channel Cembranos et al. (2017, 2022). Astrophysical data thus constrain the mass and couplings of the KK graviton states (see Fig. 3 in Ref. Cembranos et al. (2022)). In the left panel of Fig. 6 (blue area), we present the corresponding bound on $\kappa_{h}$ once the graviton mass is expressed in terms of $\rho_{T}$ and $N$. In the same panel, we display the upper bound (straight diagonal lines) due to the perturbativity constraint $N\gtrsim 5$ Agashe et al. (2007); Chen et al. (2014) for several choices of $\rho_{T}\sim\mathcal{O}(1\,\textrm{TeV})$. Since the astrophysical constraint rules out a region inside the area where the perturbativy limit is violated, the astrophysical experiments do not impose any stringent constraint on the model.
In the middle panel of Fig. 6, we present the parameter reach achieved, or expected to be achieved, at current or future colliders Cembranos et al. (2022); Aad et al. (2022, 2024). For $\rho_{T}=0.5,1,2\,$TeV, the null searches for the graviton KK modes at the LHC, for $\sqrt{s}=13$ TeV and 139 fb-1 luminosity, rule out the corresponding blue regions. At $\rho\sim\rho_{T}$, LHC imposes a strong lower bound on $N$. Still for high enough values of $N$ experimental results can be evaded as the coupling behaves as $\kappa_{h}\propto 1/N$. Analogous searches at HE-LHC and FCC-hh will expand the probed area, with a major improvement in sensitivity toward much higher values of $\rho$. In particular, for the coupling $k/M_{P}=0.1$ and a luminosity of 100 ab-1, the HE-LHC at $\sqrt{s}=27$ TeV will probe the region $m_{h}\lesssim 10$ at 95% C.L., while the LHC-hh at $\sqrt{s}=100$ TeV will probe the region $m_{h}\lesssim 100$ Helsens et al. (2019); Harris et al. (2022).
@@ VI.1 Bounds from the KK graviton sector
VI.2 Bounds from the light radion sector
The (light) radion interacts with the Higgs doublet $H=(0,v+h)^{T}/\sqrt{2}$ due to the interaction term Folgado et al. (2019)
@@ VI.2 Bounds from the light radion sector
where $m_{H}$ is the SM Higgs mass and we use the Heaviside function to swap from the low-energy to high-energy setup when $\rho$ crosses 1 TeV. Due to this interaction, the radion appears in the electroweak precision observables. We study its impact by computing its contribution to the parameters $S$ and $T$ Csaki et al. (2001); Gunion et al. (2004) and comparing it with the $\chi^{2}$ ellipse parametrization, leading to values compatible with EW observables within 95% C.L.. The outcome is presented in the right panel of Fig. 6 for different value of $\rho_{T}$ and $m_{\chi}=0.01\,\rho$ (the outcome is similar for $m_{\chi}=0.1\rho$). Due to the small coupling of the radion with the SM, the parameter space in the plane $\{\rho,N\}$ is very mildly excluded by EW observables. The exclusion mainly happens for values $\rho\sim 1\,\textrm{TeV}$, but it occurs at very small values of $N$ $(N\lesssim 5)$. This is beyond the range of reliability of the model (cf. left panel of Fig. 6), as we are neglecting gravitational quantum corrections in our treatment of the 5D model.
@@
These are the only relevant channels if we focus in the region $m_{\chi}<2m_{\mu}$, and $\Gamma_{\chi\to e\bar{e}}$ is the dominant one for $2m_{e}\lesssim m_{\chi}<2m_{\mu}$. Then the total width is $\Gamma_{\chi}\simeq\Gamma_{\chi\to\gamma\gamma}+\Gamma_{\chi\to e\bar{e}}$, and the radion lifetime $\tau_{\chi}=1/\Gamma_{\chi}$ turns out to be Koutroulis et al. (2024)
@@
A constraint follows from the fact that the radion decays would photodissociate the BBN elements if $\tau_{\chi}>\tau_{\rm BBN}\approx 10\,\textrm{s}$ Abu-Ajamieh et al. (2018), and then such a long lifetime would be forbidden. This is so except for $\tau_{\chi}>\tau_{\textrm{universe}}$, as in this case the radion can be considered stable. The lifetime that is actually excluded is then
@@
which translates into a forbidden window in the space $\{N,\rho\}$ by using Eq. (39) with some given values of $m_{\chi}$. This excluded region is shown in the left panel of Fig. 7 and rules out $2\,\textrm{MeV}\lesssim\rho\lesssim 0.1\,\textrm{GeV}$ and $10\,\textrm{MeV}\lesssim\rho\lesssim 0.3\,\textrm{GeV}$ for $m_{\chi}=0.1\rho$ and $m_{\chi}=0.01\rho$, respectively, which are natural choices for the radion mass Megías et al. (2020). The bound then varies with different choices of the model parameters that affect the radion mass.
Note that the parameter regions for which the radion lifetime fulfills the condition $\tau_{\chi}>\tau_{\rm universe}$ or barely overcomes the bound $\tau_{\chi}<\tau_{\rm BBN}$, remain questionable. In Sec. III, we compute the thermodynamic properties of the phase transition for which nucleation and percolation occur close to each other, and we can assume that reheating is practically instantaneous after percolation. A very long radion lifetime jeopardizes this approximation in the corresponding (small) region of the parameter space. Reaching firm conclusions in this regime would require more advanced techniques than those employed in this paper.
VII Summary results and conclusions
Warped extra dimensional models are among the most promising candidates for solving, or at least alleviating, the gauge hierarchy problem of the Standard Model (SM). Additionally, they naturally exhibit a strong first-order phase transition (FOPT), a key ingredient for explaining the baryon asymmetry of the universe (BAU). However, the dark matter (DM) puzzle does not allow for an easy solution: the warped embedding does not offer a stable particle meeting the DM constraints unless additional, ad-hoc ingredients are invoked Panico et al. (2008); Carena et al. (2009); Koutroulis et al. (2024). In this paper, we have demonstrated that this is not the case, as the unavoidable FOPT arising in warped models can also produce the required DM observables in the form of primordial black holes (PBHs). The satisfactory parameter window is small but appealing, as it predicts several BSM signatures that fall within the reach of multiple future experiments.
@@ VII Summary results and conclusions
Figure 8 summarizes our main results in the plane $\{\rho,N\}$, where $N$ is the number of degrees of freedom of the holographic theory, and the other parameters have been adjusted to maximize $f_{\rm PBH}$, the fraction of DM relic density in the form of PBHs that is experimentally allowed (see Fig.2). We conclude that the region with $f_{\rm PBH}=1$, which corresponds to $\rho\sim 1-1000\,$TeV, will start being probed with LVK O5, and subsequently tested by LISA and ET, which are sensitive to the SGWB generated by the radion FOPT, and later on by FCC-hh, which is sensitive to the graviton resonances. Given the multitude of BSM literature’s model proposals predicting a FOPT SGWB, the synergy between colliders and gravitational wave (GW) detectors will be of paramount importance for scrutinizing whether our warped setups are the right responsible for the phenomenology that will be observed. Finally, SGWB searches at current and future ground-based interferometers will probe the radion FOPT that maximizes the PBH abundance at $\rho\sim 10^{4}-10^{7}\,$eV. At these or higher scales, however, the warped setups lose theoretical motivation as they do not solve the hierarchy problem.
As Fig. 8 shows, multiple experiments will also test the region $\rho<1\,$TeV for which the additional intermediate brane is required. In this region, the PBH constraints forbid $f_{\rm PBH}=1$. However, the maximal abundance experimentally allowed at $\rho\sim 10^{-3}-1\,$TeV and not yet ruled out by the current LHC bound on the graviton resonances (at $\rho\sim 1\,$TeV) will be within the reach of microlensing experiments such as the Nancy Grace Roman Space Telescope (NGRST). Moreover, the narrow parameter region around our benchmark point P1 provides a good explanation of the SGWB signal hint that current pulsar timing array (PTA) experiments are finding. This region will be further probed by ET and LISA by measuring the SGWB due to the unresolved PBH population that the radion FOPT predicts. Finally, values of $\rho$ much below $10^{-4}\,$eV lead to either a radion lifetime or a FOPT entropy injection problematic for the BBN predictions.
@@
We stress that these overall conclusions are based on results assuming the radion mass to be $m_{\chi}=0.01\,\rho$, but other reasonable values of $m_{\chi}$ are not expected to substantially change our qualitative conclusions. We also remind that our estimates of the GW detector sensitivities are based on signal-to-noise ratio (SNR) computations that neglect the subtleties of the SGWB measurements, such as the existence of foregrounds, loud transient signals, and instrumental noise uncertainties. Last but not least, our estimates of the PBH abundance triggered by FOPTs rely on recent literature and are consequently evolving fast. Repeating the analysis of the present paper in more depth will surely be worthwhile when the PBH and FOPT SGWB predictions settle down.
Figure 8: Complementarity between PBH observations, the current and future GW sensitivities, and constraint/sentitivity regions from colliders. For every value of $\rho$ and $N$, $\beta/H_{\ast}$ is fixed the PBH abundance to the maximal values that current PBH constraints permit. The excluded region $\tau_{\chi}\;(\textrm{BBN})$ is displayed for $\hat{m}_{\chi}=0.1$. The bands labelled with the “SGWB” subscript assume the detection threshold SNR $>2$. The figure summarizes the results of the previous sections, which provide the details for the computation of each color band.

### v2

@@ 6 Collider and astrophysical data
The universal feature of 5D models is the presence of KK gravitons and scalar (radion) fluctuations propagating along the fifth dimension. As explained above, we are assuming the SM propagating on the TeV (or SM) brane and so the coupling of gravitons with mass $m_{h}^{n}=c_{n}\rho$ and the radion (localized toward the IR brane) depends on the distance between the SM brane and the IR brane. In such a scenario, the phenomenology is driven by the first graviton KK mode with mass $m_{h}=c_{1}\rho\equiv\hat{m}_{h}\rho$ ($\hat{m}_{h}\simeq 3.83$) and the lightest radion state with mass $m_{\chi}\equiv\hat{m}_{\chi}\rho$ ($\hat{m}_{\chi}\simeq 0.01-0.1$), which have wave functions $h_{\mu\nu}(x)$ and $\chi(x)$, respectively. [^11] We are not considering the massless graviton zero mode which is very weakly coupled to the SM as its interaction is suppressed by $1/M_{P}$. These fields are coupled to the SM with Lagrangian Csaki:2000zn
@@
$\kappa_{h}$ $=\frac{2\sqrt{2}\pi}{N}\frac{1}{\rho}\,\left[\Theta\left(\frac{\rho-\rho_{T}}{\textrm{TeV}}\right)+\left(\frac{\rho}{\rho_{T}}\right)^{2}\frac{J_{2}(c_{1}\rho/\rho_{T})}{J_{2}(c_{1})}\Theta\left(\frac{\rho_{T}-\rho}{\textrm{TeV}}\right)\right]\,,$ (34)
@@
where $J_{n}(x)$ is a Bessel function of the first kind, and $\Theta(x)$ is the Heaviside step function equal to 1 (0) for $x\geq 0$ ($x<0$) which allows us to automatically swap from the high-energy to the low-energy setup when we cross the threshold $\rho=1\,$TeV. (As explained in Sec. 2, we choose such a value of $\rho$ for concreteness, but other values of the order of 1 TeV would work as well.)
6.1 Bounds from the KK graviton sector
The existence of the KK graviton resonances influences how the strength of the gravitational force scales with distance and how fast the star’s cores lose energy due to the additional emission channel Cembranos:2017vgi ; Cembranos:2021vdv . Astrophysical data thus constrain the mass and couplings of the KK graviton states (see Fig. 3 in Ref. Cembranos:2021vdv ). In the left panel of Fig. 6 (blue area), we present the corresponding bound on $\kappa_{h}$ once the graviton mass is expressed in terms of $\rho_{T}$ and $N$. In the same panel, we display the upper bound (straight diagonal lines) due to the perturbativity constraint $N\gtrsim 5$ Agashe:2007zd ; Chen:2014oha for several choices of $\rho_{T}\sim\mathcal{O}(1\,\textrm{TeV})$. Since the astrophysical constraint rules out a region inside the area where the perturbativy limit is violated, the astrophysical experiments do not impose any stringent constraint on the model.
In the middle panel of Fig. 6, we present the parameter reach achieved, or expected to be achieved, at current or future colliders Cembranos:2021vdv ; ATLAS:2022hwc ; ATLAS:2024fdw . For $\rho_{T}=0.5,1,2\,$TeV, the null searches for the graviton KK modes at the LHC, for $\sqrt{s}=13$ TeV and 139 fb-1 luminosity, rule out the corresponding blue regions. At $\rho\sim\rho_{T}$, LHC imposes a strong lower bound on $N$. Still for high enough values of $N$ experimental results can be evaded as the coupling behaves as $\kappa_{h}\propto 1/N$. Analogous searches at HE-LHC and FCC-hh will expand the probed area, with a major improvement in sensitivity toward much higher values of $\rho$. In particular, for the coupling $k/M_{P}=0.1$ and a luminosity of 100 ab-1, the HE-LHC at $\sqrt{s}=27$ TeV will probe the region $m_{h}\lesssim 10$  TeV at 95% C.L., while the LHC-hh at $\sqrt{s}=100$ TeV will probe the region $m_{h}\lesssim 100$  TeV Helsens:2019bfw ; Harris:2022kls .
@@ 6.1 Bounds from the KK graviton sector
6.2 Bounds from the light radion sector
The (light) radion interacts with the Higgs doublet $H=(0,v+h)^{T}/\sqrt{2}$ due to the interaction term Folgado:2019sgz
@@ 6.2 Bounds from the light radion sector
where $m_{H}$ is the SM Higgs mass and we use the Heaviside function to swap from the low-energy to high-energy setup when $\rho$ crosses 1 TeV. Due to this interaction, the radion appears in the electroweak precision observables. We study its impact by computing its contribution to the parameters $S$ and $T$ Csaki:2000zn ; Gunion:2003px and comparing it with the $\chi^{2}$ ellipse parametrization, leading to values compatible with EW observables within 95% C.L.. The outcome is presented in the right panel of Fig. 6 for different value of $\rho_{T}$ and $m_{\chi}=0.01\,\rho$ (the outcome is similar for $m_{\chi}=0.1\rho$). Due to the small coupling of the radion with the SM, the parameter space in the plane $\{\rho,N\}$ is very mildly excluded by EW observables. The exclusion mainly happens for values $\rho\sim 1\,\textrm{TeV}$, but it occurs at very small values of $N$ $(N\lesssim 5)$. This is beyond the range of reliability of the model (cf. left panel of Fig. 6), as we are neglecting gravitational quantum corrections in our treatment of the 5D model.
@@
These are the only relevant channels if we focus in the region $m_{\chi}<2m_{\mu}$, and $\Gamma_{\chi\to e\bar{e}}$ is the dominant one for $2m_{e}\lesssim m_{\chi}<2m_{\mu}$. Then the total width is $\Gamma_{\chi}\simeq\Gamma_{\chi\to\gamma\gamma}+\Gamma_{\chi\to e\bar{e}}$, and the radion lifetime $\tau_{\chi}=1/\Gamma_{\chi}$ turns out to be Koutroulis:2024wjl
@@
A constraint follows from the fact that the radion decays would photodissociate the BBN elements if $\tau_{\chi}>\tau_{\rm BBN}\approx 10\,\textrm{s}$ Abu-Ajamieh:2017khi , and then such a long lifetime would be forbidden. This is so except for $\tau_{\chi}>\tau_{\textrm{universe}}$, as in this case the radion can be considered stable. The lifetime that is actually excluded is then
@@
which translates into a forbidden window in the space $\{N,\rho\}$ by using Eq. (39) with some given values of $m_{\chi}$. This excluded region is shown in the left panel of Fig. 7 and rules out $2\,\textrm{MeV}\lesssim\rho\lesssim 0.1\,\textrm{GeV}$ and $10\,\textrm{MeV}\lesssim\rho\lesssim 0.3\,\textrm{GeV}$ for $m_{\chi}=0.1\rho$ and $m_{\chi}=0.01\rho$, respectively, which are natural choices for the radion mass Megias:2020vek . The bound then varies with different choices of the model parameters that affect the radion mass.
Note that the parameter regions for which the radion lifetime fulfills the condition $\tau_{\chi}>\tau_{\rm universe}$ or barely overcomes the bound $\tau_{\chi}<\tau_{\rm BBN}$, remain questionable. In Sec. 3, we compute the thermodynamic properties of the phase transition for which nucleation and percolation occur close to each other, and we can assume that reheating is practically instantaneous after percolation. A very long radion lifetime jeopardizes this approximation in the corresponding (small) region of the parameter space. Reaching firm conclusions in this regime would require more advanced techniques than those employed in this paper.
7 Summary results and conclusions
Warped extra dimensional models are among the most promising candidates for solving, or at least alleviating, the gauge hierarchy problem of the Standard Model (SM). Additionally, they naturally exhibit a strong first-order phase transition (FOPT), a key ingredient for explaining the baryon asymmetry of the universe (BAU). However, the dark matter (DM) puzzle does not allow for an easy solution: the warped embedding does not offer a stable particle meeting the DM constraints unless additional, ad-hoc ingredients are invoked Panico:2008bx ; Carena:2009yt ; Koutroulis:2024wjl . In this paper, we have demonstrated that this is not the case, as the unavoidable FOPT arising in warped models can also produce the required DM observables in the form of primordial black holes (PBHs). The satisfactory parameter window is small but appealing, as it predicts several BSM signatures that fall within the reach of multiple future experiments.
@@ 7 Summary results and conclusions
In Figs. 8 and 9 we highlight the complementarity and synergy among PBH observations, SGWB probes, and collider tests. The two figures provide an overall view of the results detailed in the previous sections, as well as our benchmarks points P1-P8 (see Table 2). We recall that for every value of $\rho$ and $N$, the inverse FOPT duration $\beta/H_{\ast}$ is fixed such that the quantity $f_{\rm PBH}$, defined as the ratio of PBH abundance to the total dark matter abundance, reaches its experimental upper limit (see Fig. 2); much larger values of $\beta/H_{\ast}$ would relax the displayed bounds. As for the radion mass parameter $\hat{m}_{\chi}$ appearing in our collider and astrophysics computations, the choice $\hat{m}_{\chi}=0.1$ is assumed, but other reasonable values would only mildly modify the results. The FOPT SGWB sensitivity regions are based on a signal-to-noise ratio (SNR) threshold SNR $>2$, which neglects the subtleties related to the existence of foregrounds, loud transient signals, and instrumental noise uncertainties. Last but not least, our estimates of the PBH abundance triggered by FOPTs rely on recent literature and are consequently evolving fast. Repeating the analysis of the present paper in more depth will surely be worthwhile when the PBH and FOPT SGWB predictions settle down.
Bearing in mind the above caveats, Fig. 8 singles out the parameter space that is currently ruled out. The parameter region labelled as “$\tau_{\chi}\;(\textrm{BBN})$” and “$T_{R}<T_{\rm BBN}$”, corresponding to $\rho\lesssim 10^{-4}$, lead to either a radion lifetime or a FOPT entropy injection problematic for BBN and Planck observations. The region with $\rho\sim 1$ TeV conflicts with the outcome of the existing LHC searches. The LVK O3 run rules out the region $10^{4}\lesssim\rho/{\rm TeV}\lesssim 10^{7}$.
Figure 8: The regions in the plane $\{\rho,N\}$ forbidden by present SGWB searches, collider analyses on direct detection of KK gravitons, and astrophysical and cosmological observations. The bounds are based on the assumption that, at every value of $\rho$, the PBH abundance is maximized to its experimental upper limit. For other assumptions, see the sections where the bounds are derived.
On the other hand, Fig. 9 shows the overall excluded region together with the parameter reach of future colliders, gravitational-wave (GW) observatories and microlensing experiments, aiming at the detection of the KK resonances, FOPT SGWBs or PBH signatures. We conclude that the region with $f_{\rm PBH}=1$, which corresponds to $\rho\sim 1-1000\,$TeV (with the lower value in tension with LHC data), will be probed with LVK O5, and subsequently tested by LISA and ET via the FOPT SGWB observables, and later on by FCC-hh searches for graviton resonances. Given the multitude of literature’s BSM model proposals predicting a FOPT SGWB, the synergy between colliders and GW detectors will be of paramount importance for scrutinizing whether our warped setups are the right ones responsible for the phenomenology that will be observed. Finally, SGWB searches at current and future ground-based interferometers will probe the radion FOPT that maximizes the PBH abundance at $\rho\sim 10^{4}-10^{7}\,$eV. At these or higher scales, however, the warped setups partly loose theoretical motivation as they do not solve the full hierarchy problem, and the PBH abundance is orders of magnitude below the DM one (see Fig. 2).
[…]
As Fig. 9 shows, multiple experiments will also test the region $\rho<1\,$TeV for which the additional intermediate brane is required. In this region, the PBH constraints forbid $f_{\rm PBH}=1$. However, the maximal abundance experimentally allowed at $\rho\sim 10^{-3}-1\,$TeV and not yet ruled out by the current LHC bound on the graviton resonances (due to the suppressed coupling of KK gravitons to the TeV brane) will be within the reach of microlensing experiments such as the Nancy Grace Roman Space Telescope (NGRST). Moreover, the narrow parameter region around our benchmark point P1 provides a good explanation of the SGWB signal hint that current pulsar timing array (PTA) experiments are finding. This region will be further probed by ET and LISA by measuring the SGWB due to the unresolved PBH population that the radion FOPT predicts.
This landscape of both synergetic and complementary probes presents a unique opportunity for model selection, i.e. the identification of a (hopefully small) set of BSM setups compatible with the observed signatures. A generic prediction of the warped setup we have investigated is the occurrence of a highly supercooled radion-driven FOPT characterized by a large strength parameter, $\alpha\gg 1$.
For such large values of $\alpha$, the SGWB is expected to be dominated by the bubble-collision contribution, with a frequency spectrum distinct from that generated by sound waves or turbulence, which tend to dominate when $\alpha\lesssim 1$. At the same time, for $\alpha\gg 1$ and $\beta/H_{*}\lesssim 10$, which maximizes the PBH abundance, the SGWB signal is so loud that it can be precisely reconstructed provided its peak is well within an interferometer sensitivity band (see e.g. Fig. 4 in Ref. Caprini:2024hue ). Taken together, these effects suggest that if, for instance, one of our benchmarks P4-P8 is realized in nature, future GW interferometers will be able to favor setups with a large supercooling while ruling out the plethora of BSM scenarios that are either incapable of producing such a strong FOPT, or that produce it in the sound-wave or turbulance regime.
Moreover, supercooling generically implies low values of $\beta/H_{*}$, which enables a sizeble PBH production during the FOPT. In fact, as we can see from Fig. 8, a positive SGWB detection by an interferometer should be accompanied by PBH production, which could be detected by a different experiment. For instance, parameter scenarios close to benchmarks P2 and P3 would be detected by both FOPT SGWB searches at LISA and microlensing PBH imprints at NGRST. Another example is the parameter region around the benchmark P1. This region would explain the current PTA SGWB hint both in the presence and absence of a SGWB contribution due to supermassive black hole binaries. In addition, it would lead to a PBH SGWB signal detectable at ET, and PBH microlensing effects testable at NGRST.
Of course the main discriminator of the warped model compared to other BSM exhibiting strong FOPTs is the discovery of the graviton KK modes by collider experiments. For instance, the LISA discovery of the FOPT SGWB predicted around benchmarks P4 and P5 (with peak frequency $f_{p}\sim(3-6)\times 10^{-4}$ Hz and amplitude $h^{2}\bar{\Omega}_{\rm GW}\sim 10^{-8}$) would be accompanied by detection of graviton KK modes with masses $m_{\rm KK}$ in the multi-TeV range at the future HE-LHC. Similarly, the LISA discovery of a FOPT SGWB related to the parameter region between the benchmarks P6 and P7 (with peak frequency $f_{p}\sim(0.6-3)\times 10^{-3}$ Hz and amplitude $h^{2}\bar{\Omega}_{\rm GW}\sim 10^{-8}$) should be followed by discovery of graviton KK modes with masses $m_{\rm KK}\sim 10^{2}$ TeV at the future FCC-hh collider. However, for the case of the PTA SGWB detection, the graviton KK modes with masses around or below the GeV scale are too weakly coupled to ordinary matter to be tested at present or future colliders.